In right triangle ABC, AC = BC and ∠C = 90. Let P and Q be points on hypotenuse
AB, such that ∠PCQ = 45°. Show that AP^2+BQ^2=PQ^2. I think the first step is
to rotate it 90 degrees counter clockwise around C, let P go to P’ and find ∠P’CQ.

Found 3 solutions by josgarithmetic, Edwin McCravy, AnlytcPhil:
Answer by josgarithmetic(39617)

Notice the spaces added between your use of 'angle' symbol and character which follows them. Without each space, the problem description was not being fully displayed.

---
In right triangle ABC, AC = BC and < C = 90. Let P and Q be points on hypotenuse AB , such that < PCQ = 45. Show that AP^2+BQ^2=PQ^2. I think the first step is to rotate it 90 degrees counter clockwise around C, let P go to P’ and find < P’CQ.
---

Answer by Edwin McCravy(20055)
Answer by AnlytcPhil(1806)

[My previous solution had some lettering that did not match the drawn figure.  I think I have corrected all the errors below.]



Draw lines CR and RB so that ΔCRB ≅ ΔCPA
 

Now draw in QR:

 

You can finish now.  It's mostly corresponding parts of
congruent triangles and the Pythagorean theorem.  Here are 
some of the steps:

ΔCRB ≅ ΔCPA

BR = AP

∠ACP = ∠BCR  
 
∠ACP + ∠PCQ + ∠QCB = 90°

∠ACP + 45° + ∠QCB = 90°

∠ACP + ∠QCB = 45°

∠BCR + ∠QCB = 45°

Show that  ΔQCP ≅ ΔQCR

Then PQ = QR

≅QBR is a right triangle

BR² + BQ² = QR²

AP² + BQ² = PQ²

If you have trouble finishing, tell me in the space below,
and I'll get back to you be email.  (No charge, ever!!)

Edwin